Calculate the pressure in atm of 2.50 liters of CH4 gas at STP, if the volume is reduced to 2.00 liters and the temperature is raised to 227oC.
Answer #1

Given

At STP , Temperature is 273.15 K and Pressure is 1 atm.

Initial volume of methane gas ( V 1 ) = 2.50 L

Initial temperature of methane gas ( T 1) = 273.15 K

Initial pressure of methane gas ( P 1) = 1 atm

Final volume of methane gas ( V 2 ) = 2.00 L

Final temperature of methane gas ( T 2) = 500.15 K

Final pressure of methane gas ( P 2) =?

We have relation , P V = n R T

where, P is a pressure of a gas, V is a volume of a gas, n is no. of moles of gas, R is a gas constant and T is temperature of gas.

From above relation, we can write P 1 V 1 = n R T 1 (1) & P 2 V 2= n R T 2 (2)

Divide equation 2 by 1 , we get ( P 2 V 2 ) / ( P 1 V 1 ) = T 2 / T 1

( P 2 2.00 L ) / ( 1 atm 2.50 L ) = 500.15 K / 273.15 K

( P 2 2.00 L ) / ( 1 atm 2.50 L ) = 1.831

( P 2 2.00 L ) = 1.831 ( 1 atm 2.50 L )

P 2 = 1.831 ( 1 atm 2.50 L ) / 2.00 L = 2.29 atm

ANSWER : Pressure of methane gas = 2.29 atm